OpenStudy (anonymous):

Find the LCD (lowest common denominator) that eliminates the fractions in this equation. 2x/13+1/2=6

OpenStudy (anonymous):

13

OpenStudy (anonymous):

opss no. 2x/13+1/2=6 (26) 4x+13=156

OpenStudy (anonymous):

13x1=13 which is not divisible by 2. So, it is not the LCD 13x2=26 which is divisible by 2. Therefore, the ans is 26